# Curiosity-Driven Co-Development of Action and Language in Robots Through Self-Exploration

### v2

@@ Discussion
This study investigated how robots can co-develop action and language through self-exploration by integrating active inference with reinforcement learning. The experiments were designed to test four specific hypotheses, and the results provide clear support for each.
H1: Curiosity combined with motor entropy enhances developmental learning. The first hypothesis was confirmed. Robots equipped with both curiosity-driven exploration and motor entropy consistently outperformed those without, achieving higher success rates in both learned and unlearned tasks. This advantage was particularly pronounced when curiosity extended across all sensory modalities, including vision, touch, proprioception, and voice feedback. These findings suggest that the synergy of curiosity (seeking novel, unpredictable outcomes) and entropy (encouraging stochastic exploration) plays a crucial role in accelerating the acquisition of language-action mappings. This result is consistent with our previous study (?), which showed that combining curiosity and entropy significantly enhanced self-exploration in a maze navigation task. More broadly, this interpretation aligns with the active inference framework, in which agents minimize expected free energy by jointly reducing uncertainty and maximizing information gain (?, ?).
H2: Generalization follows rote learning. The results further supported the second hypothesis. In the early phases, robots were able to perform actions only for sentences encountered during training. Over time, however, they generalized to novel sentences composed of familiar words. This developmental trajectory—rote learning preceding compositional generalization—closely parallels patterns observed in human infants. Tomasello’s “verb-island” hypothesis, for example, emphasizes that children initially acquire verbs in isolated contexts before generalizing across broader structures (?). Such progression resonates with our findings, where early rigid associations gave way to flexible recombination of linguistic and motor elements.
H3: Primitive actions precede complex actions. The third hypothesis was also validated. Simpler, prerequisite-like actions such as “watch” or “be near” emerged earlier, while more complex manipulative actions like “push left” or “touch the top” developed later. This ordering mirrors hierarchical dependencies in action acquisition observed in infant development (?), reflecting how foundational skills scaffold the emergence of more sophisticated behaviors.
@@
H4: Generalization is enhanced by compositional scale. Finally, the experiments confirmed that larger vocabularies of verbs, adjectives, and nouns led to greater generalization. Robots trained with richer compositional repertoires achieved higher success rates on unlearned tasks, whereas smaller vocabularies constrained generalization severely. Our previous study (?) on supervised training of language and action for an arm robot also showed that compositional generalization improved as the size of verb–noun combinations in training increased. However, that work was limited in scale, examining only cases from $3\times 3$ to $5\times 8$ verb–noun combinations, where success rates for unlearned goals improved only from 57% to 71% under the condition of 80% training. By contrast, the present study examined much broader scaling, ranging from 48 to 180 possible compositions, while using only 33% of them for training. Under these conditions, generalization performance improved dramatically from near 25% to 90%. This contrast highlights that scale plays a critical role in enhancing generalization, and that curiosity-driven developmental learning provides a more powerful mechanism than supervised schemes under conditions of limited input. The finding also connects to the classical “poverty of the stimulus” problem raised by Chomsky (?), as it shows how compositionality enables powerful generalization from sparse training data. Once, we hypothesize that necessary training size could be proportional to summation of number of words appeared for each dimension instead of multiplication of it for all dimensions if compositionality size increases (?). This hypothesis becomes more plausible by the results in the current study which, however, should be confirmed in much more scaled experiments in the future.
@@
Taken together, these findings demonstrate that curiosity-driven exploration, motor entropy, hierarchical acquisition of actions, and scalable compositional exposure jointly support efficient co-developmental learning of language and action. The parallels with infant development—rote-to-generalization progression, prerequisite learning, and the role of vocabulary scale—suggest that the mechanisms implemented here capture essential aspects of developmental psychology. More broadly, these results strengthen the view that reconstructing developmental processes in robots can offer insights into the “poverty of the stimulus” problem, showing how powerful generalization can arise from limited input when guided by intrinsic motivation, structured experience, and the principles of predictive coding and active inference (?, ?, ?, ?).
Future Directions
@@
Future studies should extend the current framework to include interactive communication between tutors and robots. For example, when a robot cannot successfully execute a command, it could initiate a communicative act such as “Tell me how to do it” or “Ask me an easier one.” Such exchanges would allow tutors to adapt their teaching strategy dynamically, modulating the complexity of instructions or providing additional cues. This adaptive interaction resonates with Vygotskian ideas of scaffolding and the “zone of proximal development,” where caregivers adjust support according to the learner’s current abilities (?). It also aligns with research in developmental psychology emphasizing the role of joint attention, imitation, and social feedback in language learning (?, ?).
Incorporating interactive dialogue would thus move the current model closer to capturing the social nature of language acquisition in human infants. By embedding mechanisms for robots to both seek help and influence the tutoring process, future work could shed light on how social scaffolding and communicative feedback accelerate the co-development of language and action in natural developmental contexts.
Another promising future direction concerns the development of robot–robot communication through the evolution of language. Previous research has explored this possibility from different perspectives: Steels introduced the framework of “language games” to study the emergence of shared vocabularies among agents (?), Miikkulainen and colleagues investigated the evolution of artificial language through evolutionary reinforcement learning (?), and Taniguchi proposed the emergence of symbols using a collective predictive coding approach (?). While these studies have nicely demonstrated the possibility of emergent communication, they still remain limited in that they mainly achieved the emergence of object labeling or naming, whereas the evolution of action-related language, such as verbs, has been much less explored.
In this context, the current study based on active inference could be extended to address the evolution of dynamic linguistic structures, including verbs. Since our model implements active inference within a variational recurrent neural network, it is naturally suited for capturing temporal and dynamic aspects of action and language. A future extension of this work toward multi-robot interaction under the framework of “collective active inference” may thus provide novel insights into the evolution of embodied language, moving beyond static object labeling toward dynamic and action-oriented communication.
@@
This equation is derived by replacing $w$, the probabilistic model learning parameter used in Eq. Free energy principle, Active Inference, and Kawahara Model, with $z$, the probabilistic model state. The weighting coefficients $\eta$ and $\alpha$ are introduced to scale the contributions of the curiosity and entropy terms, respectively. The complexity term is computed as Kullback–Leibler divergence (KLD) between the estimated posterior distribution and the prior distribution over the latent variables at each time step. Both distributions are modeled as Gaussian distribution with time-dependent means and standard deviations. The estimated posterior is conditioned on the current sensory observation and the previous hidden state, while the prior is conditioned only on the previous hidden state. The resulting KLD thus reflects the information gain from that sensory observation, which is driven by the motor command executed at the previous time step. Therefore, exploration of more novel situations (i.e., curiosity-driven exploration) tends to result with higher information gain through larger complexity. The entropy in the third term of Eq. 3 reflects the expected uncertainty of the policy, and is computed as the negative expected log-probability of generating a motor command $a_{t}$ conditioned on the hidden state $h_{t-1}$.
@@
The forward model is trained dynamically over the course of exploratory learning by optimizing the model parameters $\psi$ to minimize the evidence free energy $F_{\psi}$ (Eq. 1) after each trial episode. The exact implementation of this process is described in the following subsection, Details of the Model Architecture.
Details of the Model Architecture
This subsection explains further details about the model architecture employed in this current study. As noted earlier, the present architecture extends our previous model (?), which is described in the “Free energy principle, Active Inference, and Kawahara Model” section of the Supplementary Materials. The primary extension involves the use of separate random latent variables, encoders, and decoders for each sensory modality. This design allows the model to process multiple types of sensation independently, including vision, tactile input, proprioception, command voice, and feedback voice. In addition, our model uses an encoder for the 4-dimensional motor command, which includes motor velocities for two the robot’s wheels and two joint angles in its arm. The full architecture of the proposed model is shown in Figure 8.
Figure 8: The details of the proposed model architecture.
Computation in this architecture proceeds as follows:
  1. 1.
  2. 2.
  3. 3.
  4. 4.
  5. 5.
  6. 6.
  7. 7.
  8. 8.
    The $Q_{t}$ value is updated according to Eq. The Employed Model.
  9. 9.
Details of the encoders and decoders of each sensory modality (e.g., vision, tactile sensation, etcetera), as well as the motor command encoder, are described in the “Implementation details” section of the Supplementary Materials.
@@ Acknowledgments
Thank you to the Okinawa Institue of Science and Technology (OIST) and colleagues in OIST’s Cognitive Neurorobotics Unit and Neural Computation Unit for supporting this work.
@@
TT was funded by OIST graduate school. J.T. was partially funded by the Japan Society for the Promotion of Science (JSPS) KAKENHI, Transformative Research Area (A): unified theory of prediction and action [24H02175],
@@
T.T. and J.T. designed the model and simulation. T.T., K.D., and J.T. designed the experiments. T.T. performed all simulation experiments and data analisus. T.T. wrote the paper. K.D. and J.T. edited the paper. J.T. supervised the study.
@@
There are no competing interests to declare.
@@
Curiosity-Driven Co-Development of Action and Language in Robots Through Self-Exploration
Theodore Jerome Tinker, Kenji Doya, Jun Tani
Cognitive Neurorobotics Unit, OIST, Onna 904-0495, Japan.
Neural Computation Unit, OIST, Onna 904-0495, Japan.
@@
$\alpha$ Importance of entropy
@@
Key5 Value10
Key5 Value10
Key5 Value10
Key5 Value10
Key5 Value10
Key5 Value10
Key5 Value10
Key5 Value10
Key5 Value10
Key5 Value10
Key5 Value10
Key5 Value10
Key5 Value10
Key5 Value10
Key5 Value10
Key5 Value10
Key5 Value10
Key5 Value10
Key5 Value10
Key5 Value10
@@
In summary, the forward model is trained to minimize the evidence free energy $F$ (Eq. S3) by accurately reconstructing sensory observations and minimizing posterior complexity based on past experiences. Meanwhile, the actor-critic pair is trained to minimize expected free energy $G$, which includes an inverted complexity term (i.e., curiosity) and entropy to encourage exploration. This leads to emergent tension in an adversarial relationship: the actor is encouraged to maximize information gain by increasing the KL divergence between prior and posterior, which the forward model trains to minimize that same term. This establishes a dynamic push-pull effect, driving self-organized exploration. Please note that the imitation term in Eq. Free energy principle, Active Inference, and Kawahara Model depends on external demonstrations or expert policies; this term is ignored in our study, which focuses on self-exploration.
@@
The robot visually senses the environment in the direction the robot faces with a $16x16x4$ image, with the four channels being red, green, blue, and distance. See Figure S1.
Figure S1: The agent’s vision, $\bm{o_{t,v}.}$ The robot is facing a magenta cone and a green pillar. The robot also sees part of its hand. The image on the left depicts the red, green, and blue channels. The image on the right depicts the distance.
@@
The second part of the sensory observation is the tactile sensation of touch. This is represented by a one value between 0 and 1 for each of the 16 sensors. Each value is equal to the fraction of time in the previous step during which the respective sensor was in contact with an object. See Figure S2.
Figure S2: The agent’s sensors for tactile sensations of touch, $\bm{o_{t,ta}}$. The robot has 16 sensors, which are planes on the surface of the robot’s body, arm, and hand. The camera and wheels are marked just for clarity.
@@
The third part of the sensation is the angle and velocity of the arm’s joints. This consists of a tensor with four values between 0 and 1: two joint angles and two joint velocities. Each value is the normalized proportion of the respective variable between its minimum and maximum range.

### v3

@@ Discussion
This study investigated how robots can develop action and language through self-exploration by integrating active inference with reinforcement learning. The experiments were designed to test four specific hypotheses, and the results provide clear support for each.
@@
H1: Generalization is enhanced by compositional scale. The experiments confirmed that larger vocabularies of verbs, adjectives, and nouns led to greater generalization. Robots trained with richer compositional repertoires achieved higher success rates on unlearned actions, whereas smaller vocabularies constrained generalization severely. Our previous study (?) on supervised training of language and action for an arm robot also showed that compositional generalization improved as the size of verb–noun combinations in training increased. However, that work was limited in scale, examining only cases from $3\times 3$ to $5\times 8$ verb–noun combinations, where success rates for unlearned goals improved only from 57% to 71% under the condition of 80% training. By contrast, the present study examined much broader scaling, ranging from 48 to 180 possible compositions, while using only 33% of them for training. Under these conditions, generalization performance improved dramatically from near 25% to 85%. This contrast highlights that scale plays a critical role in enhancing generalization, and that curiosity-driven developmental learning provides a more powerful mechanism than supervised schemes under conditions of limited input. The finding also connects to the classical “poverty of the stimulus” problem raised by Chomsky (?), as it shows how compositionality enables powerful generalization from sparse training data. Once, we hypothesize that necessary training size could be proportional to summation of number of words appeared for each dimension instead of multiplication of it for all dimensions if compositionality size increases (?). This hypothesis becomes more plausible by the results in the current study which, however, should be confirmed in much more scaled experiments in the future.
H2: Curiosity combined with motor entropy enhances developmental learning. The second hypothesis was also confirmed. Robots equipped with both curiosity-driven exploration and motor entropy consistently outperformed those without, achieving higher success rates in both learned and unlearned actions. This advantage was particularly pronounced when curiosity extended across all sensory modalities, including vision, touch, proprioception, and voice feedback. These findings suggest that the synergy of curiosity (seeking novel, unpredictable outcomes) and motor entropy (encouraging stochastic exploration) plays a crucial role in accelerating the acquisition of language-action mappings. This result is consistent with our previous study (?), which showed that combining curiosity and motor entropy significantly enhanced self-exploration in a maze navigation task. More broadly, this interpretation aligns with the active inference framework, in which agents minimize expected free energy by reducing uncertainty through maximizing information gain (?, ?).
H3: Generalization follows rote learning. The results again align with this hypothesis: in early phases, the robot succeeds only on exactly learned sentence–action pairs, and it is only over time that the robot begins to generalize to novel combinations of known elements. This mirrors developmental patterns in infants, who often begin with rigid pairings before achieving broader generalization (?). Tomasello’s “verb-island” hypothesis, for example, emphasizes that children initially acquire verbs in isolated contexts before generalizing across broader structures (?), in the same way that this robot acquired trained goals first before generalizing to untrained goals. Gerken & Knight (2015) demonstrated that 10- to 11-month-old infants can generalize from just four linguistic examples under favorable conditions (?). Moreover, Gerken et al. (2014) provide evidence that infants may generalize even from a single surprising example, suggesting that hypothesis-driven generalization can follow minimal exposure (?). These studies lend developmental credence to our observed progression from rote mapping toward flexible compositional generalization.
H4: Primitive actions precede complex actions. The fourth hypothesis was also validated. Simpler, prerequisite-like actions such as “watch” or “be near” emerged earlier, while more complex manipulative actions like “push left” or “touch the top” developed later. This ordering mirrors hierarchical dependencies in action acquisition observed in developmental psychology, where infants first master basic motor primitives before acquiring coordinated, goal-directed behaviors. Such progressive structuring of motor development has been well documented in studies showing that motor and cognitive skills emerge through iterative interaction between perception, action, and intrinsic motivation (?, ?). The dynamic systems perspective proposed by Smith and Thelen emphasizes that complex behaviors self-organize from simpler components through embodied exploration and adaptation, aligning closely with the hierarchical learning patterns observed in our robot simulations.
H5: Exception-handling rules exhibit U-shaped development. The results from Experiment 3 provide strong support for this hypothesis. When trained with two swapped command–action mappings, most robots initially produced the correct exceptional actions, subsequently regressed due to overgeneralization, and finally recovered the appropriate exception-specific mappings. This non-monotonic, U-shaped performance trajectory mirrors a well-established phenomenon in developmental psychology, in which children first succeed on irregular forms, later overgeneralize newly learned rules (e.g., producing “goed” for “went”), and ultimately reorganize their internal representations to master both rules and exceptions. Classic accounts interpret these dynamics as evidence for representational redescription—a restructuring of internal knowledge that enables more abstract, generative representations (?).
Computational modeling has long shown that such U-shaped learning can emerge naturally from error-driven or distributed representations, including Rumelhart and McClelland’s connectionist model of English past-tense acquisition (?), the multilayer perceptron models of Plunkett and Marchman (?, ?), and broader frameworks in computational developmental psychology (?). Our robot simulations demonstrate that an analogous process arises in curiosity-driven active inference: the agent first relies on rote pairings, then applies generalized compositional mappings that overwrite earlier exceptions, and finally reconstructs its latent representation to encode the exception rules correctly.
@@
Taken together, these findings demonstrate that curiosity-driven exploration, motor entropy, hierarchical acquisition of actions, and scalable compositional exposure jointly support efficient developmental learning of language and action. The parallels with infant development—rote-to-generalization progression, prerequisite learning, the role of vocabulary scale, and representation redescription—suggest that the mechanisms implemented here capture essential aspects of developmental psychology. More broadly, these results strengthen the view that reconstructing developmental processes in robots can offer insights into the “poverty of the stimulus” problem, showing how powerful generalization can arise from limited input when guided by intrinsic motivation, structured experience, and the principles of predictive coding and active inference (?, ?, ?, ?).
@@
Future studies should extend the current framework to include interactive communication between tutors and robots. For example, when a robot cannot successfully execute a command, it could initiate a communicative act such as “Tell me how to do it” or “Ask me an easier one.” Such exchanges would allow tutors to adapt their teaching strategy dynamically, modulating the complexity of instructions or providing additional cues. This adaptive interaction resonates with Vygotskian ideas of scaffolding and the “zone of proximal development,” where caregivers adjust support according to the learner’s current abilities (?). It also aligns with research in developmental psychology emphasizing the role of joint attention, imitation, and social feedback in language learning (?, ?). Incorporating interactive dialogue would thus move the current model closer to capturing the social nature of language acquisition in human infants. By embedding mechanisms for robots to both seek help and influence the tutoring process, future work could shed light on how social scaffolding and communicative feedback accelerate the development of language and action in natural developmental contexts.
Another promising future direction concerns the development of robot–robot communication through the evolution of language. Previous research has explored this possibility from different perspectives: Steels introduced the framework of “language games” to study the emergence of shared vocabularies among agents (?), Miikkulainen and colleagues investigated the evolution of artificial language through evolutionary reinforcement learning (?), and Taniguchi proposed the emergence of symbols using a collective predictive coding approach (?). While these studies have demonstrated the possibility of emergent communication in an impressive manner, they still remain limited in that they mainly achieved the emergence of object labeling or naming, whereas the evolution of action-related language, such as verbs, has been much less explored. In this context, the current study based on active inference could be extended to address the evolution of dynamic linguistic structures, including verbs. Since our model implements active inference within a variational recurrent neural network, it is naturally suited for capturing temporal and dynamic aspects of action and language. A future extension of this work toward multi-robot interaction under the framework of “collective active inference” may thus provide novel insights into the evolution of embodied language, moving beyond static object labeling toward dynamic and action-oriented communication.
@@
This equation is derived by replacing $w$, the probabilistic model learning parameter used in Eq. Free energy principle, Active Inference, and Kawahara Model, with $z$, the probabilistic model state. The weighting coefficients $\eta$ and $\alpha$ are introduced to scale the contributions of the curiosity and motor entropy terms, respectively. The complexity term is computed as Kullback–Leibler divergence (KLD) between the estimated posterior distribution and the prior distribution over the latent variables at each time step. Both distributions are modeled as Gaussian distribution with time-dependent means and standard deviations. The estimated posterior is conditioned on the current sensory observation and the previous hidden state, while the prior is conditioned only on the previous hidden state. The resulting KLD thus reflects the information gain from that sensory observation, which is driven by the motor command executed at the previous time step. Therefore, exploration of more novel situations (i.e., curiosity-driven exploration) tends to result with higher information gain through larger complexity. The motor entropy in the third term of Eq. 1 reflects the expected uncertainty of the policy, and is computed as the negative expected log-probability of generating a motor command $a_{t}$ conditioned on the hidden state $h_{t-1}$.
@@
The forward model is trained dynamically over the course of exploratory learning by optimizing the model parameters $\psi$ to minimize the evidence free energy $F_{\psi}$ (Eq. S3) after each trial episode. The exact implementation of this process is described in the supplementary material subsection, Details of the Model Architecture.
@@
Developmental learning with Exception handling rules
To examine characteristics in developmental learning with exception handling rules, we tested an alternative training setup with a pair of swapped goal-action mappings. Robots were trained such that the command “watch magenta pillar” was only rewarded if the robot performed “be near green pole” instead, and vice versa. These exceptions conflicted with the regular compositional mappings. A matched control group of identical robots was trained without such exceptions. The success-rates of these robots is depicted in Fig. 7.
@@ Acknowledgments
Thank you to the Okinawa Institute of Science and Technology (OIST) and colleagues in OIST’s Cognitive Neurorobotics Unit and Neural Computation Unit for supporting this work. We also thank Taro Toyoizumi for valuable discussions on acquiring exception handling rules.
@@
T.J.T. was funded by OIST graduate school. J.T. was partially funded by the Japan Society for the Promotion of Science (JSPS) KAKENHI, Transformative Research Area (A): unified theory of prediction and action [24H02175],
@@
T.J.T. and J.T. designed the model and simulation. T.J.T., K.D., and J.T. designed the experiments. T.J.T. performed all experiments and data analysis. T.J.T. wrote the paper. K.D. and J.T. edited the paper. J.T. supervised the study.
@@
Authors declare that they have no competing interests.
@@
Curiosity-Driven Development of Action and Language in Robots Through Self-Exploration
Theodore Jerome Tinker1, Kenji Doya1, Jun Tani1
1 Okinawa Institute of Science and Technology, Okinawa, Japan.
@@
$\alpha$ Importance of motor entropy
@@
In summary, the forward model is trained to minimize the evidence free energy $F$ (Eq. S3) by accurately reconstructing sensory observations and minimizing posterior complexity based on past experiences. Meanwhile, the actor-critic pair is trained to minimize expected free energy $G$, which includes an inverted complexity term (i.e., curiosity) and motor entropy to encourage exploration. This leads to emergent tension in an adversarial relationship: the actor is encouraged to maximize information gain by increasing the KL divergence between prior and posterior, which the forward model trains to minimize that same term. This establishes a dynamic push-pull effect, driving self-organized exploration. Please note that the imitation term in Eq. Free energy principle, Active Inference, and Kawahara Model depends on external demonstrations or expert policies; this term is ignored in our study, which focuses on self-exploration.
@@
Details of the Model Architecture
This subsection explains further details about the model architecture employed in this current study. As noted earlier, the present architecture extends our previous model (?), which is described in the “Free energy principle, Active Inference, and Kawahara Model” section of the Supplementary Materials. The primary extension involves the use of separate random latent variables, encoders, and decoders for each sensory modality. This design allows the model to process multiple types of sensation independently, including vision, tactile input, proprioception, command voice, and feedback voice. In addition, our model uses an encoder for the 4-dimensional motor command, which includes motor velocities for two the robot’s wheels and two joint angles in its arm. The full architecture of the proposed model is shown in Fig. S1.
Fig. S1: The details of the proposed model architecture.
Computation in this architecture proceeds as follows:
  1. 1.
  2. 2.
  3. 3.
  4. 4.
  5. 5.
  6. 6.
  7. 7.
  8. 8.
    The $Q_{t}$ value is updated according to Eq. The Employed Model.
  9. 9.
More thorough pseudo code is provided in the following section, utilizing variables found in table S1. Details of the encoders and decoders of each sensory modality (e.g., vision, tactile sensation, etcetera), as well as the motor command encoder, are described in the “Implementation details” section of the Supplementary Materials.
Pseudocode
The algorithm depicted in algorithm pages 1 through 3 is pseudocode representing the usage of the proposed model.
Algorithm 1 Pseudocode for Proposed Model (Part One)
Initialize forward model $f_{\psi}$, actor $\pi_{\phi}$, critic $Q_{\theta}$, replay buffer $R$ $\triangleright$ Multiple critics may be used
Initialize target critic weights $\bar{\theta}\leftarrow\theta$ $\triangleright$ One target critic for each critic
for epoch = 0, M do
$\triangleright$ In each epoch, the agent plays one episode and trains with a batch of episodes
  Initialize hidden state $h^{q}_{-1}=0.0$ and motor command $a_{-1}=0.0$ $\triangleright$ Begin new episode
  Receive observation with $n$ parts, $o_{0}=o_{0,0},...,o_{0,n}$
  for t = 0, T do $\triangleright$ Steps in episode
   for i = 0, n do $\triangleright$ Parts in observation
     $o^{enc}_{t,i}\leftarrow\textrm{enc}_{i}(o_{t,i})$ $\triangleright$ Encode observation part
[…]
$\triangleright$ Posterior inner state distribution; $||$ denotes concatenation
     $z^{q}_{t,i}\sim q(z_{t,i})=\mathcal{N}(\mu^{q}_{t,i},\sigma^{q}_{t,i}))$ $\triangleright$ Sample posterior inner state
   end for
[…]
   Execute motor command $a_{t}\leftarrow\pi_{\phi}(a_{t},h^{q}_{t})$ to receive $o_{t+1}$, $r_{t}$, and $done_{t}$
$\triangleright$ If $done_{t}$, stop episode
  end for
  Store episode’s transitions $(o_{0:T+1},a_{-1:T},r_{0:T},done_{0:T})$ in $R$ $\triangleright$ Save episode
Note: Algorithm continues.
Algorithm 2 Pseudocode for Proposed Model (Part 2)
  Sample batch of episodes $(o_{0:T+1},a_{-1:T},r_{0:T},done_{0:T},mask_{0:T})$ from $R$ $\triangleright$ Sample batch
  Initialize forward model hidden state $h^{q}_{-1}=0.0$ $\triangleright$ Begin training
  for t = 0, T+1 do $\triangleright$ Steps in advancing forward model with batch
   Initialize $P_{t-1}=0.0$ $\triangleright$ Begin tracking curiosity
   for i = 0, n do $\triangleright$ Parts in observation
     $\mu^{p}_{t,i},\sigma^{p}_{t,i}\leftarrow\textrm{MLP}^{prior}_{i}(h^{q}_{t-1}||a_{t-1})$ $\triangleright$ Prior inner state distribution
     $o^{enc}_{t,i}\leftarrow\textrm{enc}_{i}(o_{t,i})$ $\triangleright$ Encode observation part
[…]
     $P_{t-1}+=D_{KL}[q(z_{t,i})||p(z_{t,i})]$
$\triangleright$ Compare these prior and posterior to add to curiosity
     $z^{q}_{t,i}\sim q(z_{t,i})=\mathcal{N}(\mu^{q}_{t,i},\sigma^{q}_{t,i}))$ $\triangleright$ Sample posterior inner state
   end for
[…]
  end for
  for t = 0, T+1 do $\triangleright$ Steps in advancing critics with batch
   $\widehat{Q}_{t}\leftarrow Q_{\theta}(h^{q}_{t-1},a_{t})$ $\triangleright$ Predict $Q$ value
   $a^{\prime}_{t}\leftarrow\pi_{\phi}(a_{t},h^{q}_{t})$ $\triangleright$ Make new motor command with actor
   if $t>1$ then $\triangleright$ After first step, make target $Q$-values with target critics
     $\overline{Q_{t}}\leftarrow Q_{\bar{\theta}}(h^{q}_{t-1},a^{\prime}_{t})$ $\triangleright$ Get target critic’s $Q$ value
     $Q_{t}\leftarrow r_{t}+\eta P_{t}+\gamma(1-done_{t})(\overline{Q_{t}}-\alpha\mathcal{H}(\pi_{\phi}(a^{\prime}_{t}|o_{t})))$
$\triangleright$ Make target $Q$ value (eq. Free energy principle, Active Inference, and Kawahara Model)
   end if
  end for
Note: Algorithm continues.
Algorithm 3 Pseudocode for Proposed Model (Part 3)
  Initialize $F=0.0$ $\triangleright$ Initiate free energy
  for i = 0, n do $\triangleright$ Parts in observation
   $F+=(\beta_{i}D_{KL}[q(z_{0:T,i})||p(z_{0:T,i})]-\mathbb{E}_{q(z_{0:T,I})}[\log p(o_{1:T+1,i}|z_{0:T,i})])*mask_{0:T}$
$\triangleright$ Add complexity, subtract accuracy (see eq. S3)
  end for
  $\psi\leftarrow\psi-\lambda_{\psi}\frac{\partial F}{\partial\psi}$ $\triangleright$ Train forward model
  $J_{Q}(\theta)\leftarrow(\widehat{Q}_{0:T}-Q_{1:T+1})^{2}*mask_{0:T}$ $\triangleright$ Critic loss
  $\theta\leftarrow\theta-\lambda_{Q}\hat{\nabla}_{\theta}J_{Q}(\theta)$ $\triangleright$ Train critics
  $\bar{\theta}\leftarrow\tau\theta+(1-\tau)\bar{\theta}$ $\triangleright$ Update target critics
  $J_{\pi}(\phi)\leftarrow(-Q_{\theta}(o_{0:T},a^{\prime}_{0:T})-\alpha\mathcal{H}(\pi_{\phi}(a^{\prime}_{0:T}|o_{0:T})))*mask_{0:T}$
$\triangleright$ Actor loss. Use the lowest $Q$-value among critics.
  $\phi\leftarrow\phi-\lambda_{\pi}\hat{\nabla}_{\phi}J_{\pi}(\phi)$ $\triangleright$ Train actor
  if utilizing dynamic $\alpha$ with target entropy $\mathcal$$\bar${H}} then
   $J_{\alpha}(\alpha)\leftarrow\log(\alpha)\cdot(\mathcal{H}(\pi(a_{0:T}|o_{0:T}))-\bar{\mathcal{H}})*mask_{0:T}$ $\triangleright$ Alpha loss
   $\log(\alpha)\leftarrow\log(\alpha)-\lambda_{\alpha}\hat{\nabla}_{\alpha}J_{\alpha}(\alpha)$ $\triangleright$ Train alpha
  end if
end for
Note: Algorithm is completed.
@@
The robot visually senses the environment in the direction the robot faces with a $16x16x4$ image, with the four channels being red, green, blue, and distance. See Fig. S2.
Fig. S2: The agent’s vision, $\bm{o_{t,v}.}$ The robot is facing a magenta cone and a green pillar. The robot also sees part of its hand. The image on the left depicts the red, green, and blue channels. The image on the right depicts the distance.
@@
The second part of the sensory observation is the tactile sensation of touch. This is represented by a one value between 0 and 1 for each of the 16 sensors. Each value is equal to the fraction of time in the previous step during which the respective sensor was in contact with an object. See Fig. S3.
Fig. S3: The agent’s sensors for tactile sensations of touch, $\bm{o_{t,ta}}$. The robot has 16 sensors, which are planes on the surface of the robot’s body, arm, and hand. The camera and wheels are marked just for clarity.
@@
The third part of the sensation is the angle and velocity of the arm’s joints. (The velocity of the joint may not match the robot’s motor commands, because collisions with objects may restrain it.) This consists of a tensor with four values between 0 and 1: two joint angles and two joint velocities. Each value is the normalized proportion of the respective variable between its minimum and maximum range.